a lawn in the shape of a trapezoid with a heigh of 60 feet and bases of 30 and 170 feet.how many bags of fertilizer must be purchased to cover the lawn if each bag covers 2400square feet
Answer by mananth(16946)

Surface Area of a trapezium = 1/2(sum o f parallel sides) * height
A= 1/2(170+30)*60
= 6000
6000/2400 = 2.66 bags => 3 bags